OpenStudy (sleepyjess):

Proving trig equations \(\bf (sin~x)(tan~x~cos~x~-~cot~x~cos~x)~=~1~-~2~cos^2x\\\)

OpenStudy (sleepyjess):

@ganeshie8

OpenStudy (sleepyjess):

@SolomonZelman

OpenStudy (anonymous):

Start off by writing cotangent and tangent in sine and cosine terms

OpenStudy (anonymous):

focusing on the left side first

OpenStudy (anonymous):

sin(x)*tan(x)*cos(x) = sin(x)*sin(x)/cos(x)*sin(X) = sin(x)*sin(x)=sin^2(x) sinx *cot(x) *cos(x) = sin(x)*cos(x)/sin(x)*cos(x) = cos(x)*cos(X) =cos^2(x)

OpenStudy (anonymous):

Mhm

OpenStudy (sleepyjess):

tan=sin/cos cot=cos/sin

OpenStudy (anonymous):

I think we can verify it

OpenStudy (sleepyjess):

Yes I have to verify it and prove that it is true

OpenStudy (anonymous):

1= sin^2(x) + cos^(x) sin^2(x) = 1- cos^(x) 1- cos^2(x) -cos^2(x) = 1- 2*cos^2(x) this answer

OpenStudy (anonymous):

\[sinx(\frac{ sinx }{ cosx }cosx-\frac{ cosx }{ sinx }cosx)=1-2\cos^2x\]

OpenStudy (anonymous):

Yeah borak got it

OpenStudy (sleepyjess):

Ok, thank you all for the help